OpenStudy (anonymous):

multiply and divide functions f(x)=2x and g(x)=3x^3

OpenStudy (p0sitr0n):

there is only 1 function given here

Nnesha (nnesha):

to multiply or divide you should have another function g(x) or p(x)

OpenStudy (anonymous):

i fixed it im sorry

Nnesha (nnesha):

it's okay f(x) times g(x) so multiply 2x(3x^3) remember when you multiply variables you should add their exponent

OpenStudy (p0sitr0n):

ok, so function multiplication is just like any number multiplication. so you multiply constants together and x's together.

Nnesha (nnesha):

for example \[\huge\rm x \times x = x^{1+1}\]

OpenStudy (p0sitr0n):

\[x^n \times x^m= x^{n+m}\]

OpenStudy (anonymous):

so that would be 6x^3 right

Nnesha (nnesha):

nope 2x here x have exponent which is 1

OpenStudy (anonymous):

okay

Nnesha (nnesha):

so add exponent 2 times 3 = 6 right x^3 times x = ??

OpenStudy (anonymous):

^4

Nnesha (nnesha):

yep :)

OpenStudy (anonymous):

so it would be 6^4

Nnesha (nnesha):

not 6^4 x^4 so 6x^4

OpenStudy (anonymous):

okay thank you so much

Nnesha (nnesha):

np :)

Nnesha (nnesha):

don't you have to divide ?

OpenStudy (anonymous):

o yea thts the name of the page and question i think i can do the rest thanks

Nnesha (nnesha):

okay :-) gO_Od luck!! :)

OpenStudy (anonymous):

thanks if i need help is it okay if i tag you @Nnesha

Nnesha (nnesha):

sure :)

OpenStudy (anonymous):

thx

Nnesha (nnesha):

np